OpenStudy (anonymous):

(-2,2) and (0,-6) find the slope

OpenStudy (anonymous):

slope = [(y_1-y_2)\div(x_1-x_2)\] so your slope is (2--6)/(-2-0) = 8/(-2_ = -4